OpenStudy (anonymous):

Company A rents copiers for a monthly charge of $360 plus 12 cents per copy. Company B rents copiers for a monthly charge of $720 plus 6 cents per copy. What is the number of copies above which Company A's charges are the higher of the two?

OpenStudy (cwrw238):

Company A charge C = 360 + 0.12n Company B C = 720 + 0.06n

OpenStudy (cwrw238):

now you need to form an inequality where A's charges are > B's